If the position vector of one end of a chord through the focus of the parabola y^2 = 8x is 1/2i + 2j, find the position vector of the other end.


1
Expert's answer
2021-07-26T15:58:31-0400

Given parabola -


"y^{2}=8x" , the position vector of one end of an focal chord is "\dfrac{1}{2}\hat{i}+2\hat{j}"



We know that we let points on parabola as -


"(at^{2},2at)=(\dfrac{1}{2},2)"


The value of a for given parabola is -


"a=2"


So comparing the both sides of given equation , we get -


"=at^{2}=\dfrac{1}{2}"


"=t=-\dfrac{1}{2}"


Now any other point on the parabola will be "t_{1}"


Now using the property of parabola , we get -


"tt_{1}=-1"


So we get the value of "t_{1}=2" .


Now , we know that other point on focal chord of parabola is -


"(at_1^{2},2at_1)=(8,8)"


so position vector at the other end will be -


"=8\hat{i}+8\hat{j}" , which is required answer